The crab is defined as the intensity of x-rays emitted from the Crab Nebula at a given photon energy up to 30 kiloelectronvolts (keV). One crab is approximately 24 pW/m^2. What surface area in ft^2 would show the intensity of X-rays in 1 second?
1
Expert's answer
2018-10-05T14:36:09-0400
Evidently the question should be: if surface is measured in ft^2 what would be the intensity of X-rays in 1 crab?
As far as 1 foot = 0.3048 m, respectively m2 = 1/0.093 = 18.75 ft2. In this case 1 crab equals to 24/18.75 = 1.28 pW/ft^2.